### v1

# Using the antenna impedance to estimate soil electrical parameters for the MIST global 21-cm experiment

Abstract
Context. Radio experiments trying to detect the global $21$-cm signal from the early Universe are very sensitive to the electrical properties of their environment. For ground-based experiments with the antenna above the soil it is critical to characterize the effect from the soil on the sky observations. This characterization requires estimating the soil’s electrical conductivity and relative permittivity in the same frequency range as the observations.
Aims. Here we present our initial effort to estimate the conductivity and relative permittivity of the soil using the impedance of an antenna mounted at a distance above the surface. The antenna used is part of the MIST global $21$-cm experiment.
Methods. We measured the antenna impedance at three sites in the Greater Concepción area, Chile. The measurements were done between $25$ and $125$ MHz, matching the range used by MIST for sky observations. The soil parameters were estimated by fitting the impedance measurements with electromagnetic simulations of the antenna and soil. In this initial effort the soil was modeled as homogeneous.
Results. The conductivity at the three sites was found to be between $0.007$ and $0.049$ Sm-1, and the relative permittivity between $1.6$ and $12.7$. The percent precision of the estimates at $68\%$ probability is, with one exception, better (lower) than $33\%$. The best-fit simulations have a better than $10\%$ agreement with the measurements relative to the peak values of the resistance and reactance across our frequency range. For MIST, these results represent a successful proof of concept of the use of the antenna impedance for soil characterization, and are expected to significantly improve in future implementations.
Key Words.:
Astronomical instrumentation, methods and techniques – Cosmology: observations, dark ages, first stars
1 Introduction
Ground-based experiments trying to detect the sky-averaged, or global, $21$-cm signal from the early Universe are sensitive to the electrical properties of the soil at the observation site (Bradley et al. 2019; Spinelli et al. 2022; Monsalve et al. 2024a; Pattison et al. 2024). The global $21$-cm signal, produced during the Universe’s dark ages, cosmic dawn, and epoch of reionization, should be detectable at frequencies 5similar-toabsent5$\sim$ 5∼ 5200200200200 MHz (Pritchard & Loeb 2012). At these frequencies, the system temperature of radio receivers is dominated by the contribution from astrophysical foregrounds, which is at least four orders of magnitude larger than the $21$-cm signal (Helmboldt & Kassim 2009; Guzmán et al. 2011; Panchenko et al. 2013; de Gasperin et al. 2020; Sasikumar Raja et al. 2022). Due to this high system temperature, detecting the global $21$-cm signal requires the calibration of the experiment’s instrumental response, including effects from the soil, with an accuracy of at least one part in $10^{4}$.
Figure 1: MIST blade dipole antenna used to demonstrate the soil characterization technique. The antenna consists of two aluminum panels of size $1.2$ m $\times$ $60$ cm $\times$ $3$ mm, with a horizontal separation of $2$ cm. The panels are supported at a height of $50$ cm by a structure made out of fiberglass and plastic. The impedance of the antenna is measured in the gap between the two panels using an adapter. The adapter consists, at one end, of two copper wires bolted to each of the antenna panels and, at the other end, a female SMA connector. A $20$-m coaxial cable with SMA connectors is used to measure the antenna impedance with a VNA.
Several ground-based global $21$-cm experiments, such as EDGES (Bowman et al. 2018), PRIZM Philip et al. (2019), LEDA (Spinelli et al. 2021), REACH (de Lera Acedo et al. 2022), SARAS (Singh et al. 2022), MIST (Monsalve et al. 2024a), and RHINO (Bull et al. 2024) have been designed around a single, wide-beam antenna. When this type of antenna operates close to the soil relative to the observation wavelength, the electromagnetic parameters, including the impedance and beam, are directly impacted by the soil’s electrical properties (Smith-Rose & Appleton 1933; Smith-Rose 1935). To reduce the effect of the soil on the antenna performance, EDGES, PRIZM, LEDA, and REACH, have opted for operating their antennas above a metal ground plane. In the case of EDGES, PRIZM, and LEDA, the ground plane is resting on the soil, while for REACH it is lifted above the soil (de Lera Acedo et al. 2022). Although finite metal ground planes do reduce the overall sensitivity of the antenna to the soil, they could also introduce artifacts into the sky measurements that are hard to estimate and remove (Bradley et al. 2019; Mahesh et al. 2021; Spinelli et al. 2022; Rogers et al. 2022). To avoid negative effects from metal ground planes, MIST has taken the approach of conducting sky observations without a ground plane (Monsalve et al. 2024a). This choice has the trade-off of folding the full effect from the soil into MIST’s electromagnetic performance. In this context, for MIST it becomes paramount to accurately determine the soil electrical parameters and integrate them into the analysis pipelines used for calibrating the sky observations (Monsalve et al. 2024b).
In this paper, we demonstrate a technique to characterize the soil at the observation sites using the impedance of the MIST antenna. The parameters determined by this technique are the soil electrical conductivity and relative permittivity. These parameters are estimated by fitting measurements of the impedance with models from electromagnetic simulations of the antenna and soil. Our impedance measurements and simulations span the frequency range 25252525125125125125 MHz, matching the range used by MIST for observations of the sky spectrum (Monsalve et al. 2024a).
Measurements of the antenna impedance are an essential part of MIST’s calibration program. They are conducted to calibrate the impedance mismatch between the antenna and the receiver input. However, here we explore the electrical characterization of the soil, also fundamental for MIST, as an additional use of the antenna impedance. Characterizing the soil using the antenna impedance could eliminate the need for additional measurements and equipment considered for this task.
As indicated in IEEE (2011), using the impedance of an antenna located above the surface is a valid non-invasive approach for the electrical characterization of the soil at radio frequencies. This technique has been used for decades to characterize the electrical properties of materials in an antenna’s vicinity. Early results were obtained by comparing the measured impedance with analytical models (e.g., Bhattacharyya 1963; Wait 1969; Smith & King 1974; Wong et al. 1977; Nicol & Ridd 1988; He & Shen 1992). In several recent efforts the measurements have been compared with models from electromagnetic simulations (e.g., Wakita et al. 2000; Lenler-Eriksen & Meincke 2004; Berthelier et al. 2005; Le Gall et al. 2006). Our implementation of this technique represents an improvement on the latter works thanks to the availability of more powerful computers.
2 Antenna and measurements
2.1 Blade dipole antenna
Panel (a) of Fig. 1 shows the blade dipole antenna used in this work. This antenna is a replica of the antennas used by MIST for the measurements described in Monsalve et al. (2024a). The antenna consists of two solid aluminum panels of $1.2$ m in length and $60$ cm in width, mounted horizontally $50$ cm above the soil. The panel thickness is $3$ mm and the horizontal separation between the panels is $2$ cm. The antenna is supported by a structure made out of fiberglass tubes and angles, as well as plastic connectors, screws, and nuts.
The antenna is fed through an adapter located in the gap between the panels. This adapter is shown in panel (b) of Fig. 1. The adapter consists, at one end, of two copper wires bolted to each of the antenna panels and, at the other end, a female subminiature version A (SMA) connector. The antenna impedance is measured using a coaxial cable connected to the adapter. The adapter does not provide direct current (DC) isolation between the panels and the cable. Specifically, one panel has DC continuity with the center conductor of the cable and the other panel has DC continuity with the outer conductor. As it leaves the antenna, the cable runs in the middle of the gap between the panels and perpendicular to the dipole excitation axis.
The full MIST instrument also includes a receiver box installed under the antenna, which contains the electronics and batteries, and a balun connected between the antenna and the receiver box (Monsalve et al. 2024a). In this initial study, however, we test the soil characterization technique using the impedance of the antenna by itself, without the receiver box and balun.
Figure 2: Location of the test sites in the Greater Concepción area, Chile.
Table 1: Test sites in the Greater Concepción area, Chile.
Site Coordinates Surface composition
A $36^{\circ}43^{\prime}56.3^{\prime\prime}$ S, $73^{\circ}04^{\prime}14.9^{\prime\prime}$ W sand, grass
B $36^{\circ}47^{\prime}34.2^{\prime\prime}$ S, $73^{\circ}05^{\prime}45.2^{\prime\prime}$ W sand, gravel, no grass
C $36^{\circ}49^{\prime}53.5^{\prime\prime}$ S, $73^{\circ}03^{\prime}45.2^{\prime\prime}$ W clay, grass
2.2 Test sites
Measurements were conducted at three sites in the Greater Concepción area, Chile, on November 14, 2023. The sites are shown in Fig. 2 and their coordinates are listed in Table 1. Sites A, B, and C, are located on a beach, a football field, and a park, respectively. These sites were selected because they have similar characteristics to the sites used by MIST for sky observations (Monsalve et al. 2024a). At the three sites the surface is level to within a few degrees and relatively flat, with vertical variations within $\sim 5$ cm ($25$ cm) inside a $\sim 2$-m ($20$-m) radius from the antenna. Within a $\sim 20$-m radius, there are no tall plants, trees, or buildings that could cause signal reflections. At Site A, the antenna is also more than $20$ m away from the sea water. Table 1 gives a general description of the soil composition at the surface.
2.3 Measurement setup
The measurements were done with a NanoVNA-H [^1] https://nanovna.com/ vector network analyzer (VNA), in the frequency range 25252525125125125125 MHz, with a $1$-MHz resolution. The data provided by the VNA correspond to the complex reflection coefficient, $\Gamma$, which we convert to impedance, $Z$, using the following equation (Pozar 2005):
$Z=Z_{0}\frac{1+\Gamma}{1-\Gamma},$ (1)
where $Z_{0}=50$ $\Omega$ is the characteristic impedance of the measurement setup.
We connected the VNA to the antenna using a $20$-m RG174 cable with male SMA connectors. The VNA was operated manually and the operator was located by the VNA, $\sim 20$ m away from the antenna. The measurements were done with the operator sitting on the ground to minimize signal blockages and reflections. Before each antenna measurement, the setup was calibrated using the generic open, short, and load (OSL) SMA standards included with the NanoVNA-H. The calibration was done at the antenna end of the cable and, therefore, removed the unwanted S-parameters of the VNA and cable.
Figure 3: Measurements of the MIST antenna at the three test sites. (a) Reflection magnitude. (b) Reflection phase. (c) Resistance. (d) Reactance. The differences in the measurements across the sites are primarily due to differences in the electrical properties of the soils.
2.4 Measurements
The measurements are shown in Fig. 3. Panels (a) and (b) show, respectively, the magnitude and phase of the reflection coefficient. Panels (c) and (d) show, respectively, the resistance ($R$, real part of the impedance) and reactance ($X$, imaginary part of the impedance). The differences observed between the three sites are primarily due to differences in the electrical properties of the soils. The common features across sites include: (1) a dip in $|\Gamma|$ at $\sim 50$ MHz; (2) a fast transition in $\angle\Gamma$ at $\sim 50$ MHz; (3) a peak in $R$ at 65similar-toabsent65$\sim$ 65∼ 6575757575 MHz; and (4) a fast transition in $X$ at 60similar-toabsent60$\sim$ 60∼ 6080808080 MHz. Similar features are observed in the measurements done by MIST in California, Nevada, and the Canadian High Arctic (Monsalve et al. 2024a).
The measurement noise in the resistance and reactance has a sample standard deviation $<0.2$ $\Omega$, which was estimated from residuals resulting from polynomial fits to the data. For this initial study we did not estimate the systematic uncertainty in the measurements, which is primarily associated with the accuracy and stability of the OSL calibration. We expect this systematic uncertainty to dominate the measurement uncertainty budget. The determination of this uncertainty will be one of our priorities in future implementations of this technique.
Figure 4: Geometry of the Feko simulations. The simulations included the soil, $20$-m cable, antenna panels, and wires and SMA connector in the gap between the panels. The soil was simulated as homogeneous, and infinite in the horizontal directions and in depth. The meshing seen on the geometry (except on the soil) corresponds to a discretization of the surfaces for the computation of currents and fields with the method of moments.
3 Electromagnetic simulations
The electromagnetic simulations of the antenna were computed with the Feko software222www.altair.com/feko and its implementation of the method of moments. Figure 4 shows the geometry of the simulations. The geometry was designed to closely match the real antenna in order to maximize the simulation accuracy. Details of the simulation geometry are provided in Appendix A.
3.1 Soil model
The soil was simulated as homogeneous, with a perfectly flat surface, and infinite in the horizontal directions and in depth. The electrical parameters of the soil model are its conductivity ($\sigma$) and relative permittivity ($\epsilon_{r}$). This model is a first-order approximation to the real soils at our sites, in which the surface is not perfectly flat and the properties, in general, vary as a function of space. The electrical properties of real soils, in general, also vary as a function of frequency (e.g., Smith-Rose & Appleton 1933; Smith-Rose 1935; Hipp 1974; Hoekstra & Delaney 1974; Hallikainen et al. 1985; Dobson et al. 1985; Francisca & Rinaldi 2003; Bobrov et al. 2015; Sutinjo et al. 2015). In this study, $\sigma$ and $\epsilon_{r}$ were modeled as constant in the range 25252525125125125125 MHz. We made this choice under the assumption that any variations of $\sigma$ and $\epsilon_{r}$ with frequency will have a small effect on the impedance compared to our experimental uncertainties. In Feko, the effect of the soil was incorporated analytically using Sommerfeld integrals as Green’s functions for solving the boundary conditions (Davidson 2011; Mosig & Michalski 2021).
3.2 Simulation settings
The simulations were done in the range 25252525125125125125 MHz to match the impedance measurements. However, carrying out the simulations at the $1$-MHz resolution of the measurements leads to computation times of $\sim 40$ minutes per simulation on our 64-core AMD Ryzen Threadripper PRO 5995WX computer. These computation times are impractical in a pipeline that requires a rigorous sampling of the soil parameter space. The computation time roughly scales with number of frequency points requested. To decrease the computation time, we reduced the resolution of the simulations to $10$ MHz and, thus, the number of points requested from $101$ to $11$. Using $11$ complex data points evenly spaced across 25252525125125125125 MHz effectively captures the main features of the antenna impedance and is sufficient for our analysis. Other settings used in the simulations include a fine automatic mesh size and double precision calculations. With this configuration, a simulation for a single combination of soil parameters was computed in $3.5$ minutes.
4 Parameter estimation
The conductivity and relative permittivity are estimated using the simplicial homology global optimization (SHGO) algorithm available in the SciPy Python library (Virtanen et al. 2020). This algorithm, introduced in Endres et al. (2018), enables derivative-free, constrained and unconstrained optimization. SHGO minimizes the cost function using global and local searches. For the global search, SHGO incorporates the simplicial (Hatcher 2002), Sobol (Sobol 1967; Joe & Kuo 2008), and Halton (Halton 1960) sampling methods, and can also accept externally provided sampling points. In this work we use the Sobol and Halton low-discrepancy sequences for the global search. These sequences are deterministic sets of sampling points with mathematical properties comparable to, and in some cases better than, the points from the popular Latin hypercube sampling McKay et al. (1979); Renardy et al. (2021). For the local search, we use sequential least squares quadratic programming (Kraft 1988; Nocedal & Wright 2006), which is the default method in SHGO.
As described in Sec. 2.4, in this initial implementation of the technique we do not know the dominant measurement uncertainties. In the absence of uncertainty information, with SHGO we minimize the sum of squared deviations (SSD),
$\mathrm{SSD}=\sum_{i=1}^{N}{(d_{i}-m_{i})^{2}},$ (2)
where $d$ represents the data, $m$ is the impedance model from a Feko simulation, and $i$ is the index of the $N$ points available across the frequency range. At each of the $11$ frequency channels where data and model are available, the resistance and reactance are treated as two independent points and, therefore, $N=22$. The best-fit soil parameters, denoted as $\hat{\sigma}$ and $\hat{\epsilon_{r}}$, are those that produce the lowest SSD, denoted as as SSD.
For each test site and sampling sequence, we conduct a parameter estimation with SHGO in two steps. In the initial step we run SHGO with wide allowed ranges for the parameters. Specifically, we use 001111 Sm-1 for $\sigma$ and 111120202020 for $\epsilon_{r}$. These ranges include the values from all the geophysical materials that could realistically be present in the soil at our test sites (Reynolds 2011). In the final step, we run SHGO again after constraining the parameter ranges around the best-fit values found in the initial step. In both steps, we use the default settings that define the stopping criterion in SHGO. These settings lead to a number of evaluations, $N_{\mathrm{ev}}$, under $250$ in each step. Our two-step fitting approach was chosen because, starting from a wide parameter space, it can yield best-fit parameters with relatively high numerical accuracy using a total $N_{\mathrm{ev}}<500$. If we instead use a single run to explore the wide parameter space, achieving a best-fit accuracy comparable to our two-step approach would require increasing $N_{\mathrm{ev}}$ potentially by orders of magnitude.
Our fitting approach with SHGO does not directly provide uncertainties in the best-fit parameters. We determine parameter uncertainties from the SHGO samples following a separate procedure. This procedure is described in Sec. 5.3.
Table 2: Results of the parameter estimation with SHGO.
Initial step Final step
Site Sampling $\hat{\sigma}$ [Sm-1] $\hat{\epsilon}_{r}$ SSD [$\Omega^{2}$] RMSD [$\Omega$] $N_{\mathrm{ev}}$ $\hat{\sigma}$ [Sm-1] $\hat{\epsilon}_{r}$ SSD [$\Omega^{2}$] RMSD [$\Omega$] $N_{\mathrm{ev}}$
A Sobol $10^{-5}$ $1$ $1132.0$ $7.2$ $131$ $0.0070$ $1.63$ $513.3$ $4.8$ $183$
A Halton $0.0054$ $6.95$ $1880.6$ $9.2$ $103$ $0.0064$ $1.82$ $531.8$ $4.9$ $151$
B Sobol $0.0234$ $8.57$ $1869.1$ $9.2$ $131$ $0.0144$ $9.61$ $1822.6$ $9.1$ $220$
B Halton $0.0210$ $8.12$ $1861.1$ $9.2$ $103$ $0.0165$ $9.77$ $1815.5$ $9.1$ $148$
C Sobol $0.0547$ $18.66$ $930.0$ $6.5$ $137$ $0.0494$ $12.73$ $825.9$ $6.1$ $188$
C Halton $0.0523$ $16.33$ $864.2$ $6.3$ $106$ $0.0479$ $14.69$ $834.1$ $6.2$ $148$
Figure 5: Samples used in the initial step of the parameter estimation. The colors of the samples represent the SSD. The three columns correspond to the three test sites. The top (bottom) row corresponds to the Sobol (Halton) samples. In each panel, the sample with the lowest SSD, denoted as SSD, is highlighted using a red circle. In all cases, SSD occurs for $\sigma<0.06$ Sm-1.
Figure 6: Same as Fig. 5 but for the final sampling step. This step enables the identification of the best-fit values with higher accuracy than the initial step. The color scale represents different SSD ranges for the three sites.
Figure 7: Summary of the best-fit results. The top row shows the impedance measurements (solid lines), best-fit simulations (dashed lines), and best-fit soil parameter values (blue font). The bottom row shows the best-fit residuals (dashed lines) and RMSD (red font).
5 Results
5.1 Best-fit soil parameters
The results of the parameter estimation are summarized in Table 2. The table lists the best-fit soil parameters, $\hat{\sigma}$ and $\hat{\epsilon_{r}}$, and the corresponding SSD. In addition, the table shows the root mean square deviation (RMSD) between the best-fit impedance model and the data, $\mathrm{RMSD}^{\star}=\sqrt{\mathrm{SSD}^{\star}/N}$. Finally, the table shows $N_{\mathrm{ev}}$. These results are presented for the initial and final sampling steps, and for the Sobol and Halton samples.
Figure 5 shows the samples and SSDs for the initial step. The SSDs go from under 2,000 $\Omega^{2}$ to over 7,000 $\Omega^{2}$. For the three sites, the SSD decreases toward low $\sigma$, while across $\epsilon_{r}$ it remains more stable. The samples that produce the SSD are indicated using red circles. For Site A, as shown in Table 2, the best-fit values found in the initial step are $(\hat{\sigma},\hat{\epsilon_{r}})=(10^{-5}\;\mathrm{Sm}^{-1},1)$ for Sobol and $(0.0054\;\mathrm{Sm}^{-1},6.95)$ for Halton. These values are used as reference to determine adequate parameter ranges for the final step. The ranges selected for the final step are 000.020.020.020.02 Sm-1 for $\sigma$ and 111110101010 for $\epsilon_{r}$, which contain the initial best-fit values. For Sites B and C we follow a similar approach but decide to only restrict the $\sigma$ range. For these sites, the best-fit values found for $\epsilon_{r}$ in the initial step are higher than for Site A and the variation of the SSD across $\epsilon_{r}$ is low. We thus prefer to maintain 111120202020 as the range for $\epsilon_{r}$ in the final step. The $\sigma$ range is restricted to 000.040.040.040.04 Sm-1 for Site B and 000.080.080.080.08 Sm-1 for Site C.
Figure 6 shows the samples and SSDs for the final step. The final parameter spaces successfully capture the regions of lowest SSD and enable a denser sampling of these regions. The denser sampling leads to a higher accuracy in the best-fit values, which is reflected in the agreement at the $\sim 15\%$ level or better between the Sobol and Halton best-fit values. We report as the final best-fit values those that produce the lowest SSD considering both the Sobol and Halton sample sets. For Sites A, B, and C, the final best-fit values are $(\hat{\sigma},\hat{\epsilon}_{r})=(0.0070\;\mathrm{Sm}^{-1},1.63)$, $(0.0165\;\mathrm{Sm}^{-1},9.77)$, and $(0.0494\;\mathrm{Sm}^{-1},12.73)$, respectively. The best-fit values are shown in Table 3 including uncertainty ranges, which are estimated in Sec. 5.3.
5.2 Best-fit impedance simulations
Figure 7 compares the measured impedances with the best-fit simulations. The data and simulations are shown at the common $10$-MHz resolution. The simulations show good general agreement with the data, with residuals that are within ±$\pm 25$ $\Omega$ and, for most of the points, within ±$\pm 10$ $\Omega$, especially for Sites A and C. The RMSD, which considers both the resistance and reactance, is $4.8$, $9.1$, and $6.1$ $\Omega$ for Sites A, B, and C, respectively. These values represent a better than $10\%$ agreement between best-fit simulations and data when using as a reference the peak absolute values of the resistance and reactance across our frequency range, which are $>100$ $\Omega$ for the three sites.
Figure 8: Peak-normalized PDFs for $\sigma$ and $\epsilon_{r}$ obtained assuming that the uncertainty in the impedance measurements is equal to the RMSD reported in Sec. 5.2. For Site A, panel (b) shows the joint PDF for $\sigma$ and $\epsilon_{r}$, while panels (a) and (c) show the marginalized PDF for each parameter. The $68\%$ probability limits are indicated using the dashed cyan lines. We obtain $68\%$ uncertainty ranges for the best-fit parameters from the $68\%$ limits in the marginalized PDFs. Equivalent information for Sites B and C is shown in panels (d)-(i).
5.3 Uncertainties in the best-fit soil parameters
We determine uncertainties in the best-fit soil parameters reported in Sec. 5.1 using a Bayesian formalism. Bayes’ theorem indicates that the joint posterior probability density function (PDF) of the parameters given the data is
$P(\sigma,\epsilon_{r}|d)\propto\mathcal{L}(d|\sigma,\epsilon_{r})\pi(\sigma,\epsilon_{r}),$ (3)
where $\mathcal{L}(d|\sigma,\epsilon_{r})$ represents the likelihood of the data given the parameters and $\pi(\sigma,\epsilon_{r})$ represents the prior PDF. We determine the parameter uncertainties taking into account our belief that $100\%$ of the probability volume is contained within the parameter space selected for the final step of the SHGO sampling, which is shown in Fig. 6. This belief is incorporated as a uniform prior across the final parameter space.
Computing $\mathcal{L}$ typically requires knowledge of the data uncertainty. In this paper we do not know the data uncertainty and compute $\mathcal{L}$ by making some simplifying assumptions. We first assume that the uncertainty across data points is uncorrelated and for each point follows a Gaussian distribution, which leads to the following likelihood:
$\mathcal{L}\propto\exp\left[-\frac{1}{2}\sum_{i=1}^{N}{\left(\frac{d_{i}-m_{i}}{\delta_{i}}\right)^{2}}\right],$ (4)
where $\delta$ is the standard deviation of the data uncertainty. We further assume that $\delta_{i}$ is the same across data points and equal to the RMSD reported in Sec. 5.2. We make this assumption because the RMSD is the only quantitative information we have about the discrepancy between the data and a good model for the data. With this additional assumption, the likelihood becomes
$\mathcal{L}\propto\exp\left[-\frac{\mathrm{SSD}}{2(\mathrm{RMSD}^{\star})^{2}}\right],$ (5)
where $\mathcal{L}$ depends on the soil parameters only through the SSD. Equation 5 enables us to map $\mathcal{L}$, and therefore the posterior PDFs, using the same samples and SSDs produced during the SHGO parameter fitting.
The posterior PDFs of the soil parameters are shown in Fig. 8. Panels (b), (e), and (h) show the joint PDFs for Sites A, B, and C, respectively. The joint PDFs are initially sampled using the existing Sobol and Halton samples, which are shown in the panels. Joint $68\%$ probability limits were computed after interpolating the joint PDFs at high resolution between the original samples. These limits are shown using the dashed cyan lines. In Fig. 8, the other panels show the marginalized PDFs for each individual parameter. For example, panels (a) and (c) show, respectively, the marginalized PDFs for $\sigma$ and $\epsilon_{r}$ at Site A. The marginalized PDFs were computed from the high-resolution interpolation of the joint PDFs. The $68\%$ uncertainties in the soil parameters, $\delta_{\sigma}$ and $\delta_{\epsilon_{r}}$, are determined from the marginalized PDFs; specifically, from the distance between the $68\%$ limits (dashed cyan lines) and the maximum a posteriori (MAP, dashed black lines).
Our results are summarized in Table 3. For all parameters and sites, the “+” and “-” $68\%$ ranges from the marginalized PDFs are equal to within $10\%$, and we therefore report as the $68\%$ uncertainty the average of the two. The table also reports the percent precision, which for $\hat{\sigma}$ is computed as $100\%\times(\delta_{\sigma}/\hat{\sigma})$. For Site B, the percent precision of $\hat{\sigma}$ and $\hat{\epsilon}_{r}$ is, respectively, $59\%$ and $32\%$. For Sites A and C, the percent precision of $\hat{\sigma}$ and $\hat{\epsilon}_{r}$ is, respectively, $20\%$ and $25\%$.
Table 3: Best-fit values, $68\%$ uncertainties, and percent precision for the soil conductivity and relative permittivity at the three test sites.
Site $\hat{\sigma}\pm\delta_{\sigma}$ [Sm-1] $\hat{\epsilon}_{r}\pm\delta_{\epsilon_{r}}$
A $0.0070^{a}\pm 0.0014^{b}$ ($20\%$) $1.6\pm 0.4$ ($25\%$)
B $0.017\pm 0.010$ ($59\%$) $9.8\pm 3.1$ ($32\%$)
C $0.049\pm 0.010$ ($20\%$) $12.7\pm 3.2$ ($25\%$)
aBest-fit obtained as described in Sec. 5.1.
[…]
Table 4: Four-layer soil conductivity model obtained by Alegría Calvo et al. (2013) for a site $650$ m east of Site A using ERT.
Layer Thickness [m] $\hat{\sigma}$ [Sm-1]
1 (top) $0.72$ $0.00239$
2 $5.4$ $0.133$
3 $90$ $1.19$
4 (bottom) infinite $0.00011$
5.4 Comparison with expectations
In this initial effort to demonstrate the antenna impedance technique we do not attempt to offer a geophysical interpretation of our results shown in Table 3. We note, however, that our results are in broad agreement with expectations from the soil composition at the test sites. Sand and gravel, found at the surface of Sites A and B, have reported conductivity $\lesssim 0.01$ Sm-1 and relative permittivity in the range 2similar-toabsent2$\sim$ 2∼ 230303030. For clay, found at the surface of Site C, the conductivity is in the range $10^{-4}$$1$ Sm-1 and the relative permittivity in the range 2similar-toabsent2$\sim$ 2∼ 240404040. Within these ranges, the specific electrical parameter values for soils composed of sand, gravel, or clay are strongly determined by the soil moisture (Reynolds 2011).
To the best of our knowledge, the only independent soil characterization done close to our sites is the one reported in Alegría Calvo et al. (2013). They carried out electrical resistivity tomography (ERT) at a site $650$ m east of Site A. Their site is on the same beach, at coordinates $36^{\circ}44^{\prime}03.3^{\prime\prime}$ S, $73^{\circ}03^{\prime}50.4^{\prime\prime}$ W. From their measurements they obtained the four-layer conductivity model presented in Table 4.
Our conductivity estimate for Site A is not expected to match very closely the results from Alegría Calvo et al. (2013). First, being $650$ m apart, the two spots do not necessarily have similar soil characteristics. Second, while our result is valid for the range 25252525125125125125 MHz, in ERT the resistivity —inverse of conductivity— is estimated in the DC regime [^3] Electrodes are used to inject direct currents into the soil and measure the resulting voltages between different spots (Slichter 1933; Koh et al. 1997; IEEE 2011; Herring et al. 2023). Sometimes, the polarization of the injected current is reversed at frequencies $\lesssim 100$ Hz to reduce the effect of electrolytic polarization (build up of anions and cations around the electrodes) and telluric currents (currents in the ground mainly produced by the presence and fluctuation of the Earth’s magnetosphere) (IEEE 2011; Reynolds 2011; Clement et al. 2020).. Despite not expecting a good agreement, such an agreement would be mathematically possible when accounting for the spatial dependence of the results. Our result represents a weighted average of the soil conductivity over space, where the weights corresponds to the sensitivity of the antenna at different coordinates in the soil. Reporting the pattern for the antenna sensitivity in the soil falls outside the scope of this paper. However, using a reasonable set of weights for the conductivities of the four layers of the ERT model it would be possible to obtain a weighted average close to our value. A realistic situation consists of the antenna having higher sensitivity toward the top of the soil. Therefore, assigning as an example weights of $96\%$ to the top layer, $4\%$ to the second layer, and $0\%$ to the other layers, produces a weighted average of $0.0076$ Sm-1, consistent with our result $0.0070\pm 0.0014$ Sm-1.
6 Discussion and summary
Our results demonstrate that the impedance of the MIST antenna can be used for the electrical characterization of the soil in the same frequency range as the sky observations. The electromagnetic performance of MIST is very sensitive to the soil electrical properties due to MIST’s strategy of observing the sky without a metal ground plane under the antenna. Therefore, an accurate characterization of the soil at the observation sites is critical for the experiment’s success. The technique demonstrated here is particularly attractive for MIST since the measurement of the antenna impedance is already included in the calibration program of the experiment (Monsalve et al. 2024a).
The antenna impedance technique was used to estimate the soil electrical conductivity and relative permittivity at three test sites in the Greater Concepción area, Chile. The parameter estimates were determined by fitting impedance measurements with models from electromagnetic simulations of the antenna and soil. The soil was simulated as homogeneous. The measurements and simulations were conducted in the range 25252525125125125125 MHz. For the three sites, the electrical conductivity was found between $0.007$ and $0.049$ Sm-1, and the relative permittivity between $1.6$ and $12.7$. The percent precision of the estimates at $68\%$ probability is, with one exception, better (lower) than $33\%$. The best-fit simulated impedances have a better than $10\%$ agreement with the data relative to the peak absolute values of the resistance and reactance across our frequency range.
This technique has been used in the past to characterize the environment around an antenna (e.g., Wakita et al. 2000; Lenler-Eriksen & Meincke 2004; Berthelier et al. 2005; Le Gall et al. 2006) but, to our knowledge, we are the first to use it in the context of $21$-cm cosmology. Measurements of soil electrical parameters reported so far by $21$-cm experiments have been done using other techniques (Sutinjo et al. 2015; Spinelli et al. 2022). To demonstrate the impedance technique, in this initial effort we have focused on computational aspects, in particular the electromagnetic simulations and parameter estimation. We expect future refinements in the technique to enable us to reach the soil characterization accuracy required for $21$-cm cosmology.
The accuracy required in the soil parameter estimates to detect the global $21$-cm signal depends on the experiment’s design and data analysis strategy used (Pattison et al. 2024). Nonetheless, for a variety of scenarios, Spinelli et al. (2022) and Monsalve et al. (2024b) showed using signal extraction simulations that for realistic soils the accuracy required is typically better (lower) than $10\%$. The lowest parameter uncertainty reported in this paper is $20\%$ at $68\%$ probability. To characterize the soil at the sky observation sites we will use the impedance technique with the full MIST instrument instead of only the antenna panels. Using the full instrument will lead to lower uncertainties in the measurements and simulations, and therefore lower uncertainties in the soil parameters.
The impedance measurements done with the full instrument are expected to be more accurate for three main reasons. First, the instrument carries out the measurements autonomously, thus eliminating the impact of the human presence and operations. Second, the impedance is measured with a better VNA and calibrated with accurately modeled OSL standards, all of which are contained in the receiver box located under the antenna (Monsalve et al. 2024a). Third, the instrument does not use external cables or the adapter required for this paper, which introduce time instabilities that cannot be tracked and corrected.
The simulations of the full instrument are also expected to be more accurate. The cable and adapter used for this paper have a complicated geometry, which introduces inaccuracies in the simulations. These inaccuracies will be avoided in the simulations of the instrument. Further, the simulation geometry will be designed using accurate measurements of the instrument’s physical dimensions done in the field, instead of relying on nominal values and laboratory measurements as in this paper.
In addition to increasing the accuracy in the measurements and simulations, quantifying their uncertainties represents a priority for future implementations of this technique. Robust uncertainty estimates will enable us to select between competing models for the soil as we attempt to make the model more realistic, and determine reliable uncertainties for the soil parameters.
This soil characterization technique is directly relevant for global $21$-cm experiments planning to observe the sky from the surface of the lunar farside (e.g., Bale et al. 2023). The electrical characteristics of the lunar regolith vary significantly with location (Feng et al. 2020; Siegler et al. 2020). Therefore, to reach the calibration accuracy required for $21$-cm cosmology, the conductivity and relative permittivity of the regolith will have to be determined at each observation site. Measurements of the antenna impedance primarily conducted for radiometer calibration could also be used for in-situ regolith characterization.
Acknowledgements.
We thank Roberto Vives and Benjamín Fernández for their help with the antenna construction and measurements. We also thank Matheus Pessôa for his work during early tests of the antenna impedance technique. We acknowledge support from ANID Chile Fondo 2018 QUIMAL/180003, Fondo 2020 ALMA/ASTRO20-0075, Fondo 2021 QUIMAL/ASTRO21-0053, and Fondo 2022 ALMA/31220012. We acknowledge support from FONDECYT Fondo Iniciación 11221231. We acknowledge support from Universidad Católica de la Santísima Concepción Fondo UCSC BIP-106. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2019-04506, RGPNS 534549-19. We acknowledge the support of the Canadian Space Agency (CSA) [21FAMCGB15]. This research was undertaken, in part, thanks to funding from the Canada 150 Research Chairs Program. This research was enabled in part by support provided by SciNet and the Digital Research Alliance of Canada.
Appendix A Geometry of the Feko simulations
Here we provide details of the Feko simulation geometry, which is shown in Fig. 4. The geometry was designed based on nominal dimensions and laboratory measurements of the antenna assembly. This assembly was replicated at the three test sites as closely as possible.
In the simulated geometry, the surface of the soil is perfectly flat and aligned with the $xy$-plane at $z=0$. The antenna panels are simulated as aluminum boxes of $1.2$ m in length, $60$ cm in width, $3$ mm in thickness, and $2$ cm in horizontal separation. The length and width of the panels are aligned parallel to the $xy$-plane at a height $z=50$ cm. The long dimension of the panels is aligned with the $x$-axis. The gap between the panels is centered at $(x,y,z)=(0,0,50\;\mathrm{cm})$.
The wires of the adapter in the gap between the panels are simulated as copper cylinders with a diameter of $0.64$ mm. The wires make contact with the panels while aligned with the $x$-axis at the center of the gap. As they leave the panels, the wires bend in the $+y$ direction, where they extend for $6.3$ cm before reaching the SMA connector. The connector is simulated as a gold cuboid of $1.6$ cm in length and $6$ mm in width and thickness. The electrical excitation of the antenna is applied by a voltage port located at the junction between one of the wires and the SMA connector. The location of the port mimics the impedance calibration point for the real antenna at the SMA connector.
The cable is simulated as a $2$-mm diameter copper cylinder leaving the antenna in the $+y$ direction. This cylinder begins at the output of the SMA connector and runs along the middle of the gap between the panels up to the edge of the panels. At the edge of the panels, the cylinder bends downward $18.3^{\circ}$ and continues in diagonal until it reaches a lowest height of $5$ mm above the soil. This point occurs at a distance of $1.5$ m from the edge of the panels in the $+y$ direction. The cylinder then starts running parallel to the soil in the $+y$ direction at a height of $5$ mm until it reaches a total length of $20$ m.
In the simulations we use the following conductivities (Pozar 2005): for aluminum, $3.816\times 10^{7}$ Sm-1; for copper, $5.813\times 10^{7}$ Sm-1; and for gold, $4.098\times 10^{7}$ Sm-1. The support structure of the antenna, made out of fiberglass and plastic, does not affect the antenna impedance and is not included in the simulations. Other features not included are the panel mounting holes. These holes would significantly increase the complexity of the geometry and computation time but according to our tests do not have a noticeable effect on the impedance.